OpenStudy (anonymous):

What is the reciprocal of the fraction -3/7? Is it 7/3 or -7/3? I'm not completely sure...

OpenStudy (anonymous):

7/-3? -7/3, i'd think

OpenStudy (anonymous):

I know what a reciprocal is, but I'm not sure if the fraction is supposed to be negative or positive...

OpenStudy (across):

The reciprocal of a number \(a\) is a number \(b\) such that \(ab=1\).

OpenStudy (anonymous):

@ilikequadraticsokay they are the same thing -7/3

OpenStudy (anonymous):

yeah I was just thinking to myself lol sorry

OpenStudy (across):

Apparently, you don't know what a reciprocal is. ;) Let \(a=-3/7\) and let \(b=7/3\). Then \(ab=(-3/7)(7/3)=-1\). Does this meet our definition?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

I still don't understand...

OpenStudy (anonymous):

So is it -7/3?